Math 210 F1 S98 — Calculus III — Home Quiz #1 — Hanson

Show your own work on this sheet in about 15 minutes (10pts/ques). Sign your name to testify that this is your own work: Name . SocSecNo. .

1. Describe the region of R^{3} represented by x^{2} + z^{2} ≤ 2 and 0 ≤ y ≤ 3, including a sketch.

2. Find |a⃗|, |b⃗|, and a⃗ − 2b⃗ if a⃗ =< 3, −4, 5 > and b⃗ =< 2, 3, −1 >.

3. Find a⃗ • b⃗, if a⃗ =< 3, −4, 5 > and b⃗ =< 2, 3, −1 >.

4. Find a⃗ × b⃗, if a⃗ =< 3, −4, 5 > and b⃗ =< 2, 3, −1 >.